OpenStudy (anonymous):

Hi, may I ask, I have y'+(x^n)y=0. Given x=1, ln(y)=-1/4 and x=2, ln(y)=-4. I need to find n. May I know when we perform variable separable on DE, do we include constant C?

OpenStudy (kc_kennylau):

what's DE?

OpenStudy (anonymous):

Differential equation?

OpenStudy (abb0t):

If you weren't given initial conditions, you do need to include C. But "C" is just a constant.

OpenStudy (abb0t):

But, if I am reading your question correctly, you do have initial conditions, so you do need to solve for constant, C.

OpenStudy (anonymous):

I know, but the problem is, I do not know n so how can I solve C?

OpenStudy (kc_kennylau):

Use the initial conditions

OpenStudy (kc_kennylau):

One for solving n and one for solving c

OpenStudy (anonymous):

oh? we are gonna use it separately? but i actually thought of doing it simultaneous equation way, and cancel out C in order to find n but I was stuck half way

OpenStudy (anonymous):

use the initial condition

OpenStudy (kc_kennylau):

Bear in mind that x=1, ln(y)=-1/4 is ONE condition

OpenStudy (anonymous):

Yes, that is true, but then the way to solve is it integrate differential eq and then substitute the conditions in it to get 1st eq right? Then i think we will end up with n and C. Do we do the same and get a 2nd equation using the 2nd condition then do simultaneous eq?

OpenStudy (anonymous):

\[-\ln |y|= \frac{ x^{n+1} }{ n+1 }+C\]

OpenStudy (anonymous):

the above eq is what i got from the integration of Diff Eq. So, is it right?